OpenStudy (anonymous):

solve the differential equation y''+y'-3y=0.

myininaya (myininaya):

\[r^2+r-3=0 => r=\frac{-1 \pm \sqrt{1 - 4(-3)}}{2}=\frac{-1 \pm \sqrt{13}}{2}\] this implies \[y=c_1 e^{\frac{-1+ \sqrt{13}}{2}}+c_2e^\frac{-1 - \sqrt{13}}{2}\]

myininaya (myininaya):

oops missing something

OpenStudy (amistre64):

yeah, its sqrty

myininaya (myininaya):

\[y=c_1 e^{\frac{-1+ \sqrt{13}}{2}x}+c_2e^{\frac{-1 - \sqrt{13}}{2}x}\]

myininaya (myininaya):

i was missing the little x

OpenStudy (anonymous):

I know I can count on someone here Thank you all.

myininaya (myininaya):

:)